OpenStudy (arcane):

How much work does the electric field do in moving a -7.7μC charge from ground to a point whose potential is +65 V higher?

OpenStudy (irishboy123):

The potential energy of the charge in the field is \(U = q V\) So the **change** in PE is \(\Delta U = q \Delta V = (-7.7 \times 10^{-6} C)(65V)\) The work actually done by the field is \(W = - \Delta U\)

OpenStudy (irishboy123):

Or the better way IMHO, but depending upon what you know, is to take the given: \(E = - \dfrac{dV}{dx}\) And then to note that the work done by any force is: \(W = \int F(x) ~ dx\) Which in the case of an E field is: \(= q\int E(x) ~ dx\) \(= - q\int \dfrac{dV}{dx} ~ dx\) \(\implies W = - q \Delta V\)